OpenStudy (rippedblue42):

What is the sum of a 7-term geometric series if the first term is 6, the last term is 24,576, and the common ratio is 4? Plz halp!!!

OpenStudy (rippedblue42):

Plz help!!!

OpenStudy (loser66):

I forgot what geometric series is :( Can you state it out?

OpenStudy (rippedblue42):

What do you mean?

OpenStudy (legomyego180):

Is this calc 2? @loser66 A geometric series is a series in the form: \[\sum_{k=a}^{\infty}(\frac{ a }{ b })^k\] Where if \(\frac{ a }{ b } > 1\) the series diverges, \(\frac{ a }{ b } < 1\) converges. The sum formula for a geometric series is: \[\sum_{k=c}^{\infty}(\frac{ a }{ b })^k \rightarrow _{SUM}=\frac{ (\frac{ a }{ b } )^c}{ 1-\frac{ a }{ b } }\]

OpenStudy (sshayer):

\[S _{7}=6 \times \frac{ 4^7-1 }{ 7-1 }=?\]

OpenStudy (sshayer):

sum of a geometric series with c.r r>1 is \[S _{n}=a \times \frac{ r^n-1 }{ r-1 }\] where a is first term ,n=number of terms r=common ratio.

OpenStudy (rippedblue42):

It's algebra 2

OpenStudy (rippedblue42):

Is it 19,662?

OpenStudy (rippedblue42):

Idk what do to with the common ratio

OpenStudy (rippedblue42):

???

OpenStudy (rippedblue42):

Oh wait. It'd be 32,766

OpenStudy (loser66):

@legomyego180 thanks for the explanation :) I wish I get the answer from the Asker :)

OpenStudy (faiqraees):

\[\large\rm S_n=\frac{initial(1-ratio^{terms})}{1-ratio} \]